The half-life of uranium-235 is much longer than the half-life of carbon-14,

so it can be used to determine the age of rocks that are too old to be dated

with carbon dating.

A scientist finds a fossilised reptile encased in rock containing 82% of the

normal amount of uranium-235. Use the half-life graph above to determine

approximately how long ago the reptile lived.


1
Expert's answer
2021-09-02T02:01:14-0400



t=0.639/k
700=0.639/k
k=0.693/700
=9.9*10^-4
t (82/100)=0.1985/9.9*10^-4

=200.5 million years